The question is:
How much must you invest each quarter in a bond fund yielding 9.7% compounded quarterly to become a millionaire in 20 years?
For some reason I'm completely stuck.

Answer by edjones(7311)

Let A=1,000,000, r=.097, n=4, t=20 P=principal
.
A=P(1+ r/n)^nt
1000000=P(1+ .097/4)^(4*20)
P=10^6/(1.02425)^100
=10^6/10.9799...
=$91,075.75 Final answer
.
Ed